OpenStudy (anonymous):

Solve the following triangles. If the answer isnt a whole number, round to the hundreths place 1. A=60 degrees, a=1, b=2 I NEED HELP!!

OpenStudy (anonymous):

I dont see a diagram, but I'm assuming A is the angle in between a and b. So use Area = 0.5 x a x b x Sin 60

OpenStudy (anonymous):

do you need the area?

OpenStudy (anonymous):

no I just have to find angle B, Angle C and side c

OpenStudy (anonymous):

Use the Sine and Cosine rules

OpenStudy (anonymous):

Law of sines is what we are doing

OpenStudy (campbell_st):

the law of sine says \[\frac{\sin(60)}{1} = \frac{\sin(B)}{2}\] solve for B then use angle sum of a triangle to find C then use the law of sine again \[\frac{1}{\sin(60)} = \frac{c}{\sin(C)}\] hope it helps